# Active particle motion in Poiseuille flow through rectangular channels

### v1

@@ V.4.4 Poincaré map and sticky trajectories
Since there exists one constant of motion $H_{s}$ for our $4$D dynamical system in Eq. (9), the effective dynamics of the system take place in $3$D. We can further explore the nature of the system dynamics and transition to chaos by using Poincaré sections which allows us to visualize regular and chaotic regions of the system on a $2$D plot. We obtain a Poincaré map by sampling the active particle trajectory at times $t_{n}$ that correspond to the intersection of the phase-space trajectory with the $\theta=0$ (equivalently $e_{y}=0$) hyperplane, i.e. $\theta(t_{n})=0$ with $\dot{\theta}(t_{n})>0$. At these times $t_{n}$, we store the values $\phi_{n}=\phi(t_{n})$ and $x_{n}=x(t_{n})$, and plot them against each other giving us a Poincaré map. We repeat this for active particle trajectories with different initial conditions but for a fixed value of the constant of motion $H_{s}$ i.e. the initial conditions satisfy
$H_{s}=$ $\frac{1}{2}\bar{U}\left(x(0)^{2}+y(0)^{2}-x^{2}(0)y^{2}(0)\right)$ (15)
@@
In this way, we obtain Poincaré maps as shown in Fig. 7 for different values of the constant of motion $H_{s}$. We find that for small values of $H_{s}$, we observe regular behavior with ubiquitous quasiperiodic trajectories that correspond to closed curves on the Poincaré map (see Fig. 7(a)-(b)). As $H_{s}$ is increased, we observe that, due to nonlinear resonances, some of these orbits break into a chain of smaller orbits (see Fig. 7(c)). Further increase in $H_{s}$ gives rise to chaos as evident by the apparently random scatter of points in the Poincaré map (see Fig. 7(d)-(e)). At these large values of $H_{s}$, we have a mixture of order and chaos where small islands of regular behaviors exist within the chaotic sea. Increase in the value of $H_{s}$ can also be interpreted as initial positions going away from the center of the channel. For fixed upstream initial orientation $\theta(0)=0$, $\phi(0)=0$, Eq. (15) describes a closed curve
$x(0)^{2}+y(0)^{2}-x^{2}(0)y^{2}(0)=2H_{s}/\bar{U}.$
@@
These closed curves are the same shape as the contours of the flow field $\bar{u}(x,y)$. Small values of $H_{s}$ correspond to circle-like closed curves near the center of the channel while increasing $H_{s}$ leads to square-like closed curves away from the center of the channel. Since near the center of the channel, the dynamics of the system are similar to a circular cross-section and hence integrable, this transition to chaos with increasing $H_{s}$ maybe understood in terms of the theory of nearly integrable Hamiltonian systems and KAM theory [27].
@@
When islands of regular behaviors exist within the chaotic sea of a Poincaré map, “sticky” trajectories can arise where a long time is spend in the vicinity of these periodic islands [28]. An example of such a sticky tumbling trajectory is shown in Fig. 8 and supplemental video S11. In this plot, panels (a) and (b) show the trajectory in different planes while panel (c) shows a Poincaré map with the light green colored part in all three panels representing the “sticky” behavior. Such trajectories spend a very long time near the boundaries of the periodic islands compared to the time spent in a domain of chaotic sea of the same phase-space volume.
V.4.5 Dynamics in the large $\bar{U}$ limit
@@
Figure 9: Active particle dynamics for large $\bar{U}=50$ starting at $z(0)=0$ with fixed initial upstream orientation angles $\theta(0)=0$ and $\phi(0)=0$. (a) A typical confined trajectory in the $(z,y)$ (left) and $(x,y)$ (middle) plane, as well as the orientation evolution (right) for initial position $x(0)=0.6,y(0)=0.1$. (b) Classification of behaviors in the $(x(0),y(0))$ initial position plane using the color scheme in Fig. 2. Poincaré maps in the $(x_{n},\phi_{n})$ plane for $\theta(t_{n})=0$ with $\dot{\theta}(t_{n})>0$ for the constant of motion (c) $H_{s}=1.75$ and (d) $H_{s}=3$. These values of constant of motion are indicated by black curves in panel (b). The trajectory in panel (a) was simulated for $t=100$ in the same units as Eq. (9), which corresponds to $t=100\,\bar{U}$ in the units of Eq. (16). See also supplemental videos S12 for a video of the active particle trajectory in panel (a).
@@
In the limit of large $\bar{U}\gg 1$, Eq. (9) approach a singular limit for the evolution of $\theta$ and $\phi$. Hence, to understand this regime, we rescale the dimensionless time by $\bar{U}$ in Eq. (9) (with $AR=1$) and obtain the following system:
$\dot{x}$ $=-\frac{1}{\bar{U}}\cos\theta\sin\phi$ (16)
$\dot{y}$ $=\frac{1}{\bar{U}}\sin\theta$
@@
This form of the system removes the singular terms and allows efficient numerical solution. A typical classification of trajectories at a large value of $\bar{U}=50$ along with an example trajectory and Poincaré maps is shown in Fig. 9. In this regime of large $\bar{U}$ we find that chaos ceases and we have regular behaviors. A typical regular trajectory that confines itself to a section of the channel is shown in Fig. 9(a). Simulating many different initial conditions with fixed initial upstream orientation $\theta(0)=0$ and $\phi(0)=0$ gives us the classification in the initial position plane $(x(0),y(0))$ as shown in Fig. 9(b). Compared to Fig. 6, we find that the central green region still exists while the cyan, blue and red regions have almost vanished. Outside the central region, the behavior is dominated by off-centered trapped trajectories. Two different Poincaré maps at $H_{s}=1.75$ and $H_{s}=3$ are shown in Figs. 9(c) and (d) respectively. These typically show closed loops indicating quasiperiodic behavior of the system in this regime. With variation in the initial orientation angles $\theta(0)$ and $\phi(0)$ from the upstream equilibrium state, we find that in the initial position plane, some yellow regions transition to purple regions corresponding to regular tumbling trajectories.
Hence, from Figs. 6 and 9 we see that the active particle motion is regular for small and large $\bar{U}$, whereas chaos emerges for intermediate $\bar{U}$.
@@ V.5 Active particle transport along the channel
Figure 10: Transport of the active particle along the straight channel based on the initial position in the $(x(0),y(0))$ plane for various $\bar{U}$ and fixed $\theta(0)=0$ and $\phi(0)=0$. Final position of the active particle in the $z$ direction starting at $z(0)=0$ (color) for values of (a) $\bar{U}=0.5$, (b) $\bar{U}=2.5$, (c) $\bar{U}=5$, (d) $\bar{U}=7.5$, (e) $\bar{U}=10$ and (f) $\bar{U}=15$.
The trajectories shown in Fig. 2 started at $z=0$ and the time series is shown for $t=900$ to $t=1000$. We see that the transport of the active particle along the $z$ direction i.e. along with the flow, can vary significantly depending on the type of active particle motion realized in the channel cross-section. Active particles that perform swinging motion near the center of the channel (e.g. green, blue, cyan and some red trajectories) will travel further along the channel compared to active particles whose motion is confined near the edges of the cross-section (e.g. purple trajectory). Figure 10 shows for different $\bar{U}$ values, a contour plot of the $z$ location at the end of the simulation ($t=1000$) in the $(x(0),y(0))$ plane for initially upstream-oriented active particles (starting from $z=0$). For a small value of $\bar{U}=0.5$, we see that the final $z$ positions are negative indicating that the active particle’s intrinsic speed dominates the fluid flow speed resulting in a net upstream migration of the active particle. For a larger value of $\bar{U}=2.5$, the fluid speed dominates particle speed and we obtain a net downstream migration of all active particles. Further increase to $\bar{U}=5$ does not change the transport profile near the center of the cross-section, however, near the corners, we see fluctuations in the transport profile due to the appearance of chaotic wandering motion in this region (see Fig. 6(c)). For larger values of $\bar{U}$ the active particle transport is dominated by the background fluid flow profile. We find a central plug region corresponding to large transport of the active particles undergoing swinging motion near the center of the channel where flow speed is large, while near the edges and corners we observe small transport of active particle corresponding to off-centered trapping or tumbling trajectories that stay near the outer regions of the cross-section where the flow speed is small. In between the large transport central plug region and the small transport outer region, we observe some fluctuations in the transport contours that correspond to chaotic wandering trajectories.
@@ VI Dynamics in wider rectangular channels
In this section, we briefly explore the effects of aspect ratio $AR$ of the rectangular cross-section on the active particle dynamics. To understand this, we plot the trajectory classification and LLE in the initial condition plane $(x(0),y(0))$ for two different rectangular cross-sections with $AR=2$ and $AR=4$ as shown in Fig. 11(a) and (b) respectively, while keeping $\theta(0)=0$, $\phi(0)=0$ and $\bar{U}=10$ fixed. We note that the classification criteria presented for square channels in Sec. V.1 has been scaled based on $AR$ in the $x$ direction for rectangular channels. So, for example, the classification of central swinging motion (green) has been modified to confinement in a rectangular box of domain $-0.25\,AR<x<0.25\,AR$ and $-0.25<y<0.25$. For the rectangular cross-section with $AR=2$, we find similar types of active particle trajectories at similar initial positions compared to the $AR=1$ square channel (see Fig. 4) with some minor differences. Near the center of the channel, we obtain central swinging motion (green) as well as vertical (cyan) and horizontal (blue) swinging motion. We do see a horizontal stretching of these regions due to widening of the channel. Moreover, the relative frequency of oscillations in the horizontal and vertical direction near the center of the channel will be scaled by $AR$ as per the eigenvalues in Sec. III.0.1. Beyond the central region, we find wandering motion (red) and tumbling motion (purple) similar to the square channel. We find that the yellow islands, corresponding to confined trajectories away from the channel center, have diminished in size compared to the square cross-section, and these regions are more scattered. A typical trajectory in the yellow region is shown in Fig. 11(c) where the motion is trapped in an off-centered vertical band compared to the motion confined near the corner for a square cross-section (see Fig. 2(d)). For the $AR=2$ channel, we also see the emergence of small horizontal swinging (blue) regions appearing near the left and right edges of the horizontal centerline that were not present for square channels. The LLE also shows similar structure with regular trajectories near the central region of the channel and the dominance of chaos in the outer regions of the channel. The $AR=4$ rectangular cross-section shows similar features as the $AR=2$ channel with a few noteworthy differences. The yellow regions for off-centered trapping motion increases in extent while the small regions for horizontal swinging present for $AR=2$ channel no longer exist. Moreover, the tumbling motion (purple) region penetrates the wandering motion (red) region near the left and right edges of the cross-section. We also note that our simple flow field approximation in Eq. (8) will become poor near the left and right edges of the cross-section with increasing $AR$, and hence a more accurate flow field for Poiseuille flow in rectangular channels may be needed to accurately capture the active particle dynamics in these regions.
@@
Figure 11: Active particle trajectory classification (top) and LLE (bottom) in the initial position plane $(x(0),y(0))$ for motion in rectangular channels with (a) $AR=2$ and (b) $AR=4$. (c) A trajectory of a typical yellow region with $x(0)=0.75$ and $y(0)=0.55$ for $AR=2$ channel. Other parameters are fixed to $\bar{U}=10$, $z(0)=0$, $\theta(0)=0$ and $\phi(0)=0$. See supplemental video S13 for a video of the active particle trajectory in panel (c).
@@ VII Conclusions
In summary, we identified a general constant of motion for a point-like active particle interacting with a steady unidirectional fluid flow. We studied in detail, theoretically and numerically, the specific case of an active particle in Poiseuille flow through straight channels with rectangular cross-section. The equations of motion resulted in a $4$D nonlinear dynamical system with one constant of motion. We identified two equilibrium states for this particle-fluid system located at the center of the rectangular cross-section: (i) an upstream-oriented marginally stable equilibrium where small perturbations lead to oscillatory motion about this equilibrium point and (ii) a downstream-oriented unstable saddle equilibrium. By numerically solving the system, we observed a variety of active particle trajectories for different values of $\bar{U}$ and $AR$, as well as different initial conditions. These were classified based on the regions they occupy in the channel cross-section. Swinging trajectories such as central swinging, vertical swinging and horizontal swinging are typically observed as quasiperiodic motion about the center of the channel, where as off-centered trapping motion typically occurs as confined quasiperiodic motion away from the channel centerline. Tumbling trajectories stay near the edges of the cross-section while wandering trajectories visited both the central and the outer regions of the cross-section. By calculating the largest Lyapunov exponents, many of the tumbling and wandering trajectories were shown to be chaotic. Poincaré maps with increasing value of the constant of motion showed the transition to chaotic behavior and the persistence of small islands of regular behavior in the chaotic sea. This resulted in “stickiness” for chaotic tumbling trajectories where the chaotic trajectory gets trapped near a periodic state for a long time.
@@
We studied how the active particle motion varies with the system parameters and initial conditions. With changing aspect ratio $AR$ of the channel, we obtain the same qualitative types of trajectories. For variations with $\bar{U}$, we observe rich dynamics with non-chaotic motion at very small and large $\bar{U}$ whereas chaos emerges in an intermediate range of $\bar{U}$. In this regime of $\bar{U}$ where chaos arises, we found that the active particle trajectories are generally very sensitive to initial conditions with a couple of robust regimes. The active particle oriented upstream and starting near the channel center typically undergoes regular swinging motion that is robust with small variations in the initial position and orientation. Similarly, the active particle oriented upstream and starting near the edges of the channel typically undergoes chaotic tumbling motion which is again robust with small variations of the initial conditions.
@@
The model used in this paper is simple and can be extended in various ways to more accurately capture the motion of natural and artificial microswimmers in channel flows. The present model does not capture the interaction of the active particle with the channel walls and it would be useful to explore the effects of wall interactions to (i) understand motion of active particle in narrow channels and (ii) more accurately captures the active particle trajectories get too close to the channel walls. The present work highlights the important of initial conditions on active particle motion and hence it would be useful to perform careful microfluidic experiments with active particle to quantify the effects of initial conditions on the active particle motion in channel flows. By considering wall interactions using squirmer models for the active particle in $2$D planar Poiseuille flow, Zöttl and Stark [16] showed that adding wall interactions results in emergence of dissipative dynamical features such as a stable point attractor and limit cycle attractor for the upstream-orientation swinging motion. Choudhary et al. [20] explored the effects of adding fluid inertia for active particles in $2$D planar Poiseuille flow and reported similar dynamical features. Our previous work on the dynamics of passive spheres in $3$D channel flows with non-zero fluid inertia has revealed rich dynamical structure for inertial particle focusing behaviors [29, 30, 31]. In future work, we aim to understand the effects of inertia on the dynamics and focusing of active particle in $3$D microfluidic channel flows.

### v2

@@ V.4.4 Poincaré map and sticky trajectories
Since one constant of motion $H_{s}$ remains for our $4$D dynamical system in Eq. (8), the effective dynamics of the system take place in $3$D. We can further explore the nature of the system dynamics and transition to chaos by using Poincaré sections to visualize regular and chaotic regions of the system on a $2$D plot. We construct a Poincaré map by sampling the active particle trajectory at times $t_{n}$ that correspond to a crossing of the phase-space trajectory with the $\theta=0$ (equivalently $e_{y}=0$) hyperplane in the positive direction, i.e. $\theta_{n}=\theta(t_{n})=0$ with $\dot{\theta}_{n}=\dot{\theta}(t_{n})>0$. At these times $t_{n}$, we store the values $\phi_{n}=\phi(t_{n})$ and $x_{n}=x(t_{n})$, and plot them against each other giving us a Poincaré map. We repeat this for many active particle trajectories having different initial conditions but a common fixed value of the constant of motion $H_{s}$, i.e. the initial conditions satisfy
$H_{s}=$ $\frac{1}{2}{U}\left(x(0)^{2}+y(0)^{2}-x^{2}(0)y^{2}(0)\right)$ (13)
@@
Several such Poincaré maps are shown in Fig. 7 for different values of the constant of motion $H_{s}$. We find that for small values of $H_{s}$, we observe regular behavior with ubiquitous quasiperiodic trajectories that correspond to closed curves on the Poincaré map (see Fig. 7(a,b)). As $H_{s}$ is increased, we observe that, due to nonlinear resonances, some of these orbits break into a chain of smaller orbits (see Fig. 7(c)). Further increase in $H_{s}$ gives rise to chaos as evident by the apparently random scatter of points in the Poincaré map (see Fig. 7(d,e)). At these large values of $H_{s}$, we have a mixture of order and chaos where small islands of regular behaviors exist within the chaotic sea. Increase in the value of $H_{s}$ can also be interpreted as initial positions going away from the center of the channel. For example, with fixed initial upstream orientation $\theta(0)=\phi(0)=0$, Eq. (13) describes a closed curve
$x(0)^{2}+y(0)^{2}-x^{2}(0)y^{2}(0)=2H_{s}/{U}.$
@@
These closed curves are the same shape as the level sets of the flow field ${u}(x,y)$. Small values of $H_{s}$ correspond to circle-like closed curves near the center of the channel while increasing $H_{s}$ leads to square-like closed curves away from the center of the channel. Since near the center of the channel, the dynamics of the system are similar to a circular cross-section and hence integrable, this transition to chaos with increasing $H_{s}$ may be understood in terms of the theory of nearly integrable Hamiltonian systems and KAM theory [27].
@@
When islands of regular behaviors exist within the chaotic sea of a Poincaré map, “sticky” trajectories can arise where a long time is spend in the vicinity of these periodic islands [28]. An example of such a sticky tumbling trajectory is shown in Fig. 8 and supplemental video S11. In this plot, panels (a) and (b) show the trajectory in different planes while panel (c) shows a Poincaré map with the light green colored part in all three panels representing the “sticky” behavior. On the Poincaré map, such trajectories spend a very long time near the boundaries of the periodic islands compared to the time spent in a domain of the chaotic sea of the same phase-space volume.
V.4.5 Dynamics in the large ${U}$ limit
@@
Figure 9: Active particle dynamics for large ${U}=50$ starting at $z(0)=0$ with fixed initial upstream orientation angles $\theta(0)=0$ and $\phi(0)=0$. (a) A typical confined trajectory in the $(z,y)$ (left) and $(x,y)$ (middle) planes, as well as the orientation evolution (right) for initial position $x(0)=0.6,\ y(0)=0.1$. (b) Classification of behaviors in the $(x(0),y(0))$ initial-position plane using the color scheme of Fig. 2. (c,d) Poincaré maps in the $(x_{n},\phi_{n})$ plane for $\theta_{n}=0$ with $\dot{\theta}_{n}>0$ for the constants of motion (c) $H_{s}=1.75$ and (d) $H_{s}=3$. These values of the constant of motion are indicated by black curves in panel (b). The trajectory in panel (a) was simulated for $t=100$, using the time scaling of Eq. (8), which corresponds to $t=100\,{U}$ using the time scaling of Eq. (14). See also supplemental videos S12 for a video of the active particle trajectory in panel (a).
@@
In the limit of large ${U}\gg 1$, Eq. (8) approaches a singular limit for the evolution of $\theta$ and $\phi$. Hence, to understand this regime, we rescale the dimensionless time in Eq. (8) (with $AR=1$) by ${U}$ to obtain the following system:
$\dot{x}$ $=-\frac{1}{{U}}\cos\theta\sin\phi$ (14)
$\dot{y}$ $=\frac{1}{{U}}\sin\theta$
@@
This form of the system removes the singular terms and allows efficient numerical solution. A typical classification of trajectories at a large value of ${U}=50$ along with an example trajectory and Poincaré maps are shown in Fig. 9. In this regime of large ${U}$ we find that chaos ceases and we have regular behaviors. A typical regular trajectory that confines itself to a quadrant of the channel is shown in Fig. 9(a). Simulating many different initial conditions with fixed initial upstream orientation $\theta(0)=\phi(0)=0$ gives us the classification in the initial position plane $(x(0),y(0))$ as shown in Fig. 9(b). Compared to Fig. 6, we find that the central green region still exists while the cyan, blue and red regions have almost vanished. Outside the central region, the behavior is dominated by off-centered trapped trajectories (yellow region). Two different Poincaré maps at $H_{s}=1.75$ and $H_{s}=3$ are shown in Figs. 9(c) and (d) respectively. These typically show closed loops indicating quasiperiodic behavior of the system in this regime. With variation in the initial orientation angles $\theta(0)$ and $\phi(0)$ from the upstream equilibrium state, we find that in the initial-position plane, some yellow regions transition to purple regions corresponding to regular tumbling trajectories.
Hence, from Figs. 6 and 9 we see that the active particle motion is regular for small and large ${U}$, whereas chaos emerges for intermediate ${U}$.
@@ V.5 Active particle transport along the channel
Figure 10: Axial transport of the active particle along the straight channel based on the initial position in the $(x(0),y(0))$ plane for various ${U}$ and fixed $\theta(0)=0$ and $\phi(0)=0$. The colour bars indicate the final axial coordinate $z$ of the active particle from an initial axial position of $z(0)=0$ for values of (a) ${U}=0.5$, (b) ${U}=2.5$, (c) ${U}=5$, (d) ${U}=7.5$, (e) ${U}=10$ and (f) ${U}=15$.
The trajectories shown in Fig. 2 started at $z=0$ and the time series is shown for $t=900$ to $t=1000$. We see that the transport of the active particle in the $z$ direction i.e. axially along with the flow, can vary significantly depending on the type of active particle motion realized in the channel cross-section. Active particles that perform swinging motion near the center of the channel (e.g. green, blue, cyan and some red trajectories) will travel further along the channel compared to active particles whose motion is confined near the walls of the channel (e.g. purple trajectories). Figure 10 shows, for different ${U}$ values, a contour plot of the $z$ (axial) location at the end of the simulation ($t=1000$) in the $(x(0),y(0))$ plane for initially upstream-oriented active particles (starting from $z=0$). For a small value of ${U}=0.5$, we see that the final axial positions are negative indicating that the active particle’s intrinsic speed dominates the fluid flow speed resulting in a net upstream migration of the active particle. For a larger value of ${U}=2.5$, the fluid speed dominates the particle speed and we obtain a net downstream migration. Further increase to ${U}=5$ does not qualitatively change the axial transport profile near the center of the cross-section, however, near the corners, we see fluctuations in this profile due to the appearance of chaotic wandering motion in this region (see Fig. 6(c)). For even larger values of ${U}$ the active particle axial transport is dominated by the background fluid flow profile. We find a central plug region corresponding to large axial transport of the active particle undergoing swinging motion near the center of the channel where flow speed is large, while near the walls and corners we observe small axial transport corresponding to off-centered trapping or tumbling trajectories that stay near the outer regions of the cross-section where the flow speed is small.
@@ VI Dynamics in wider rectangular channels
In this section, we briefly explore the effects of the aspect ratio $AR$ of the rectangular cross-section on the active particle dynamics. In Fig. 11(a,b) we plot the trajectory classification and LLE in the initial-position plane $(x(0),y(0))$, while keeping $\theta(0)=0$, $\phi(0)=0$ and ${U}=10$ fixed, for two different rectangular cross-sections with $AR=2$ and $AR=4$, respectively. We note that the classification criteria presented for square channels in Sec. V.1 has been scaled based on $AR$ in the $x$ direction for rectangular channels. So, for example, the classification of central swinging motion (green) has been modified to confinement in a rectangular box of domain $-0.25\,AR<x<0.25\,AR$ and $-0.25<y<0.25$. For the rectangular cross-section with $AR=2$, we find similar types of active particle trajectories at similar initial positions compared to the $AR=1$ square channel (see Fig. 4) with some minor differences. Near the center of the channel, we obtain central swinging motion (green) as well as vertical (cyan) and horizontal (blue) swinging motion, but there is a horizontal stretching of these regions due to increasing the width of the channel. Moreover, the relative frequency of oscillations in the horizontal and vertical directions near the center of the channel will be scaled by $AR$ as per the eigenvalues in Sec. III.0.1. Beyond the central region, we find wandering motion (red) and tumbling motion (purple) similar to the square channel. We find that the yellow islands, corresponding to confined trajectories away from the channel center, have diminished in size compared to the square cross-section, and these regions are more scattered. A typical trajectory in the yellow region is shown in Fig. 11(c) where the motion is trapped in an off-centered vertical band compared to the motion confined near the corner for a square cross-section (see Fig. 2(d)). For the $AR=2$ channel, we also see the emergence of small horizontal swinging (blue) regions appearing near the left and right ends of the horizontal centerline that were not present for square channels. The LLE also shows similar structure with regular trajectories near the central region of the channel and the dominance of chaos near the channel walls. The $AR=4$ rectangular cross-section shows similar features to the $AR=2$ channel with a few noteworthy differences. The yellow regions for off-centered trapping motion increase in extent while the small regions for horizontal swinging present for $AR=2$ channel no longer exist. Moreover, the tumbling motion (purple) region penetrates the wandering motion (red) region near the left and right edges of the cross-section. We also note that our simple flow field approximation in Eq. (7) will become poor near the left and right edges of the cross-section as $AR$ increases, and hence a more accurate flow field for Poiseuille flow in rectangular channels may be needed to accurately capture the active particle dynamics in these regions.
@@
Figure 11: (a,b) Active particle trajectory classification (top) and LLE (bottom) in the initial position plane $(x(0),y(0))$ for motion in rectangular channels with (a) $AR=2$ and (b) $AR=4$. (c) A typical trajectory of a particle starting in the yellow region with $x(0)=0.75$ and $y(0)=0.55$ for the case $AR=2$. Other parameters are fixed to ${U}=10$, $z(0)=0$, $\theta(0)=0$ and $\phi(0)=0$. See supplemental video S13 for a video of the active particle trajectory in panel (c).
@@ VII Conclusions
We have studied in detail, theoretically and numerically, the motion of a point-like active particle in a steady unidirectional fluid flow, specifically through a straight channel with rectangular cross-section. We identified a general constant of motion that enabled the six equations of motion to be reduced to a $4$D nonlinear dynamical system with one constant of motion. We identified two equilibrium states for this particle-fluid system located at the center of the rectangular cross-section: (i) an upstream-oriented marginally stable equilibrium where small perturbations lead to oscillatory motion about this equilibrium point and (ii) a downstream-oriented unstable saddle equilibrium. By numerically solving the system, we observed a variety of active particle trajectories for different values of the maximum flow speed ${U}$ and the channel width/height aspect ratio $AR$, as well as different initial particle positions and orientations. The trajectories were classified based on the regions they occupy in the channel cross-section. Swinging trajectories, such as central swinging, vertical swinging and horizontal swinging, were typical quasiperiodic motions near the centerlines of the channel, whereas off-centered trapping motion was the typical form of confined quasiperiodic motion away from the channel centerlines. Tumbling trajectories stay near the walls of the channel while wandering trajectories visited both the central and the outer regions of the cross-section. By calculating the largest Lyapunov exponents, many of the tumbling and wandering trajectories were shown to be chaotic. Poincaré maps with increasing value of the constant of motion showed the transition to chaotic behavior and the persistence of small islands of regular behaviour in the chaotic sea. The latter resulted in “sticky” chaotic tumbling trajectories due to the chaotic trajectory becoming trapped near a periodic state for a long time.
@@
We have shown how the active particle motion varies with the system parameters and initial conditions. We focused on a square channel cross-section ($AR=1$) and also showed that qualitatively similar particle trajectories were obtained in cross-sections with larger aspect ratio $AR$. Varying the maximum flow speed ${U}$, revealed rich dynamics with non-chaotic motion at very small and large ${U}$, and the emergence of chaos in an intermediate range of ${U}$. In this regime of ${U}$ where chaos arises, we found that the active particle trajectories are generally very sensitive to initial conditions with a couple of robust regimes. The active particle oriented upstream and starting near the channel center typically undergoes regular swinging motion that is robust to small variations in the initial position and orientation. Similarly, the active particle oriented upstream and starting near the walls of the channel typically undergoes chaotic tumbling motion which is again robust to small variations of the initial conditions.
@@
The model used in this paper is simple and can be extended in various ways to more accurately capture the motion of natural and artificial microswimmers in channel flows. The present model does not capture the interaction of the active particle with the channel walls and it would be useful to explore the effects of wall interactions to (i) understand motion of active particles in narrow channels and (ii) more accurately capture the active particle trajectories that get very close to the channel walls. The present work highlights the importance of initial conditions on active particle motion and hence it would be useful to perform careful microfluidic experiments to quantify the effects of initial conditions on active particle motion in channel flows. By considering wall interactions using squirmer models for the active particle in $2$D planar Poiseuille flow, Zöttl and Stark [16] showed the emergence of dissipative dynamical features such as a stable point attractor and a limit cycle attractor for the upstream-orientation swinging motion. Choudhary et al. [20] explored the effects of adding fluid inertia for active particles in $2$D planar Poiseuille flow and reported similar dynamical features. Our previous work on the dynamics of passive spheres in $3$D channel flows with non-zero fluid inertia has revealed rich dynamical structure for inertial particle focusing behaviors [29, 30, 31]. In future work, we aim to understand the effects of inertia on the dynamics and focusing of active particle in $3$D microfluidic channel flows.